last month we had 18 days of rain or snow. There were 8 more days of snow than rain. How many days did it snow?
Answer by ikleyn(52781)

.

    Snow    + Rain = 18

    S       +  R   = 18

    (R + 8) +  R   = 18

    2R + 8         = 18

    2R             = 18 - 8 = 10

     R                      = 10/2 = 5.


ANSWER.  5 days Rain;  5 + 8 = 13 days Snow.

I completed my solution.

Is everything clear to you ?

If you still have questions, do not hesitate to post them to me.